# Point-hyperplane incidences via extremal graph theory

### v1

@@ 6 Concluding remarks
In this paper we proved tight bounds on the maximum number of incidences between $n$ points and $m$ hyperplanes in a $d$-dimensional space $\mathbb{F}^{d}$, assuming their incidence graph is $K_{s,s}$-free. More precisely, for every $(m,n)$, we can find a field $\mathbb{F}=\mathbb{F}(d,m,n)$ for which our upper bounds can be realized up to a constant factor depending only on $d$ and $s$, for sufficiently large $s\geq s_{0}(d)$. It would be interesting to improve our upper bounds for a fixed finite field $\mathbb{F}_{p}$ and arbitrary $m$ and $n$. In two dimensions such estimates were obtained by Bourgain, Katz, and Tao [4].
Another interesting problem is to understand the size of the largest bipartite graph in dense incidence graphs. Given a set of $n$ points $\mathcal{P}$ and a set of $m$ hyperplanes $\mathcal{H}$ in $\mathbb{R}^{d}$ with $I(\mathcal{P},\mathcal{H})\geq\frac{mn}{2}$, what are the optimal bounds on the minimum of $\operatorname{rs}(\mathcal{P},\mathcal{H})$? Singer and Sudan [25] (based on unpublished works of Lovett and Pálvölgyi and Fox, Wigderson) proved that there is a configuration of $n$ points $\mathcal{P}$ and $n$ linear hyperplanes $\mathcal{H}$ such that the points and the normal vectors of the hyperplanes have only 0-1 coordinates, $I(\mathcal{P},\mathcal{H})\geq n^{2}/2$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})=n^{2}2^{-\Theta(\sqrt{d})}$. Lovett [21] conjectured that this bound is the best possible, using the equivalent terminology of low-rank matrices. This problem might be also interesting over fields other than $\mathbb{R}$. For every field $\mathbb{F}$ the lower bound $\operatorname{rs}(\mathcal{P},\mathcal{H})=\Omega(\frac{mn}{2^{2d}d})$ follows from a recent result of Singer and Sudan [25]. On the other hand, by taking all points and all hyperplanes in $\mathbb{F}_{2}^{d}$, we have $n=2^{d}$, $m=2(2^{d}-1)$, $I=\frac{mn}{2}$, and it is not difficult to show that every complete bipartite graph in the incidence graph has at most $2^{d}\approx\frac{mn}{2^{d+1}}$ edges.

### v2

@@ 7 Concluding remarks
In this paper, we proposed a novel combinatorial approach to study incidence problems over arbitrary fields. We considered several natural problems, such as determining the maximum number of point-hyperplane incidences, point-variety incidences, and unit distances, under standard non-degeneracy assumptions.
Our bounds are also optimal in the sense that they cannot be improved unless assumptions on the underlying field are made. Hence, it would be interesting to decide whether the techniques in this paper can be combined with some field-dependent information to improve some of our bounds. Such improved bounds are known over the real numbers [2, 22], and also in some cases over finite fields [5, 43]. For example, Rudnev’s theorem [43] shows that when $n\leq p^{2}$, there are at most $O(m\sqrt{n}+ms)$ incidences between $m$ points and $n$ planes in $\mathbb{F}_{p}^{3}$, assuming the incidence graph is $K_{2,s}$-free.
Finally, it would be also interesting to find other geometric problems for which the methods of this paper can be applied. For example, it may be possible to use our approach to address other non-degeneracy conditions beyond forbidding $K_{s,s}$ in the incidence graph.
Acknowledgements. We would like to thank Larry Guth for suggesting to study point-variety incidences and stimulating discussions. We would also like to thank Joshua Zahl and Misha Rudnev for pointing out additional references.